OpenStudy (anonymous):

Explain how to use the graph of f(x) = 2x to obtain the graph of g(x) = log2x.

OpenStudy (eyust707):

hmm im kinda curious...

OpenStudy (anonymous):

do you mean f(x)=2 to power x or just f(x)=2x

OpenStudy (anonymous):

Explain how to use the graph of f(x) = 2^x to obtain the graph of g(x) = log2x. sorry about the confusion.!

OpenStudy (anonymous):

now I know how to answer your question first, you need to know the relationship between exponential function and log function. they are the inverse function of each other

OpenStudy (anonymous):

for example , the point (1,2) is on the graph of f(x)=2^x, then the point (2,1) will be on the graph of g(x)=log2x. Also, their graphs is symmetric by y=x

OpenStudy (anonymous):

Ok thanks, I totally understand what you are saying.